e2:

Which step could be used to help isolate the variable in the following equation? 5.6j – 0.12 = 4 + 1.1j

e2:

i need help

Hero:

subtract 1.1j from both sides then add 0.12 to both sides: 

e2:

so it going to be 4.5j=4.12?

e2:

and so i divide it?

e2:

Hero:

Okay, which of those options do you think is the correct one. Think carefully before answering.

e2:

is it 1.1j?

Hero:

State the entire statement or choose A,B, C, or D.

e2:

i press D

Hero:

Option D is Correct

e2:

thank you

e2:

Solve the equation. –3x + 1 + 10x = x + 4 A)x = x equals StartFraction one-half EndFraction B)x = x equals StartFraction 5 Over 6 EndFraction C)x = 12 D)x = 18

Hero:

Use operations (addition/subtraction) to group like terms on the same side